A body falling freely under gravity from a tower of height h covers a distance of 9/16h during the last second of its motion.then the height of tower is ???

Grade:11

1 Answers

sonika p
6 years ago
h = 0.5* g * t^2
t = root (2h/g)
distance travelled in last second, 9/16h = 0.5g(2h/g) – 0.5g(root(2h/g) – 1)
simplifying we get, 81 h2 -368gh +64g2 = 0
h = 4.36m, 0.18m